# Swarm-intelligent search for gravitational waves from eccentric binary mergers

### v5

@@ II.3 Multiple detector optimization
When searching in the coincident mode, the processes described earlier in section II.2 can, in principle, proceed independently and in parallel for all the available detectors. However, in such a case, the templates are also computed independently for each detector and the point of convergence can differ due to several reasons. When projected from arbirary directions in the sky, the antenna pattern functions of the detectors modulate the incoming signal differently, thus effectively shifting the expected optimal point. In addition, the noise contributions from the individual detectors can slightly shift the convergence point. Otherwise, several local peaks form an elongated band of almost degenerate templates with similar SNRs as shown in Fig. 3, which can confuse the convergence. But a true gravitational wave signal should have the times of arrival and the source parameters consistent across all the available detectors. To ensure the second criterion, the same set of template parameters should optimize a fitness function computed for the detector network. Besides, for a given template, a stretch of data can contain several noise triggers, however, such triggers are expected to be uncorrelated in time and appear at just one detector. A data stretch here represents the data recorded by the available detectors within a fixed GPS start and end times. Thus, to minimize spurious optimization of glitch triggers, a natural choice is to jointly optimize the reweighted SNRs across the network of detectors. The joint fitness function for a network of detectors is given by,
@@
The steps involved in the optimization process is described as follows. These are also summarized in Fig. 4. To demonstrate the process, we assume that the coincident strain data from different detectors exist in the computer storage. The data blocks for each detector are divided into segments overlapping with their immediate neighbours in time to account for the corrupted ends of the SNR timeseries due to filtering artifacts. The duration of the data stretch is chosen in a way that an in-band inspiral signal lies completely within a data stretch. The PSO algorithm explores each data stretch one after another. Data segments are downsampled till 1024 Hz as discussed later in Section III. PSDs used in the matched-filters are computed from the segments using Welch’s method (Welch, 1967).
@@
Based on the updated values of the fitness function of each particle, the p-best and the g-best locations and fitnesses are computed as described in subsection II.2. Following this, Eq. 2-3 are used to assign a new set of positions to the particles for the next iteration. The algorithm also checks whether any particle has drifted beyond the desired search boundaries- a reflecting boundary puts back an outlier inside in a symmetric position while flipping its velocity 111While this is suitable in most cases, we find that an adsorbing, boundary is more effective for the eccentric dimension towards the recovery of non-eccentric sources, where instead of putting the outlier back to a symmetric location, we put it on the boundary.. The algorithm additionally checks for a terminal condition of a convergence below a small mean radius of the swarm or a preset timeout, whichever is earlier. The boundary conditions can improve the termination and also ensures no hang-up occurs due to non-availability of template parameters outside the region allowed by a waveform model. Note that these steps, namely evolving the position-velocity, applying the boundary conditions and the termination checks are computationally trivial. At this point, the current iteration is complete.
@@ II.4 Significance of candidates
When the iterations finish, the algorithm returns the detection template for a given stretch of data with the associated triggers from the individual detectors. The detection template point is already consistent in the source parameters across all the detectors. Now the secondary triggers are tested for any zero-lag time coincidence to create foreground events. A time coincidence test requires triggers from different detectors to occur within a time window defined by the light’s travel time between the detectors plus 5 ms accounting for any possible timing measurement errors. Several such foreground candidates, which pass both the parameter and the time coincidence tests, may still not be of astrophysical origin. To assess the possibility of a noise origin, one assigns a statistical significance to a foreground candidate and/or a probability of being of astrophysical origin, (Farr et al., 2015; Capano et al., 2017). Here we assign a false alarm rate (FAR) to the foreground events using the method of time-shifting of triggers obtained from processing the data immediately available near (or before) the candidate (Wąs et al., 2009). The algorithm maintains a sufficient buffer of the cumulative secondary triggers resulting from the detection templates of the neighbouring (or past) data stretches. Note that the different stretches can have their optimized g-best template points located throughout the parameter space. Thus, though the secondary triggers for a given data stretch have consistency over the source parameters, the cumulative of such triggers over multiple stretches do not enforce the source parameter consistency 222An appropriate strategy to utilize even the primary triggers for the background estimates in the PSO search can be explored in a future study. However, the cumulative triggers from the entire space are generally more prone to picking up abrupt noise and can even degrade the background statistic.. If the triggers obtained from a pair of detectors are artificially shifted with respect to each other by a time more than travel time of light between the two detectors, then the resulting coincidences cannot be of astrophysical origin. Such artificial background events are also ranked by their artificial network statistic value and a FAR estimate is assigned to a foreground candidate as
@@ III.3 Background event rates
In the previous sections, we have shown that the eccentric searches can outperform the aligned-spin searches in the presence of Gaussian noise. However, the data from a real detector can contain nearby noise glitches which can overwhelm the searches. To investigate this, we start with an hypothesis that there is no strong astrophysical signal and data segments contain only detector noise and glitches. We perform the searches independently using the same configuration as described in section III on several data blocks from the third observing run (O3a) of the Advanced LIGOs not containing any obvious CBC candidates (Abbott et al., 2023, 2021e). For simplicity we use data from only two detectors H1 and L1. Each 4096 seconds data block is divided into segments of 256 seconds with a 50% overlap, equally shared with its neighbouring segments and the analysis procedure described in Sec. II.3 is followed. Thus each PSO instance effectively searches 128 seconds of data from the available detectors. Note that the triggers crossing an SNR threshold undergo the same signal consitency tests and the network reweighted SNR is used as the ranking statistic. For comparison, we also conduct the aligned-spin non-eccentric search using PSO on the same data segments with identical set-up. The results are summarized in Fig. 11.
@@
However, to assess the accuracy of the estimation of the false alarm rates described earlier in subsection II.4, we use the triggers obtained from analyzing the above datasets. The cumulative number of all the foreground candidates are plotted against a given inverse false alarm rate (IFAR) threshold in Fig. 12. Here the FARs are assigned using the cumulative triggers from both BNS and BBH searches, inclusive of those forming the foreground candidates. The expected number of noise events for a given FAR threshold over a coincident observation time is calculated using Eq. 6. The figure demonstrates a general self consistency of the FARs assigned to any of the foreground candidates and indicates that the candidates are consistent with noise origin in this case.

### v6

@@ II.3 Multiple detector optimization
When searching in the coincident mode, the processes described earlier in section II.2 can, in principle, proceed independently and in parallel for all the available detectors. However, in such a case, the templates are also computed independently for each detector and the point of convergence can differ due to several reasons. The noise contributions from the individual detectors can slightly shift the convergence point. Otherwise, several local peaks form an elongated band of almost degenerate templates with similar SNRs as shown in Fig. 3, which can confuse the convergence. But a true gravitational wave signal should have the times of arrival and the source parameters consistent across all the available detectors. To ensure the second criterion, the same set of template parameters should optimize a fitness function computed for the detector network. Besides, for a given template, a stretch of data can contain several noise triggers, however, such triggers are expected to be uncorrelated in time and appear at just one detector. A data stretch here represents the data recorded by the available detectors within a fixed GPS start and end times. Thus, to minimize spurious optimization of glitch triggers, a natural choice is to jointly optimize the reweighted SNRs across the network of detectors. The joint fitness function for a network of detectors is given by,
@@
The steps involved in the optimization process are described as follows. These are also summarized in Fig. 4. To demonstrate the process, we assume that the coincident strain data from different detectors exist in the computer storage. The data blocks for each detector are divided into segments overlapping with their immediate neighbours in time to account for the corrupted ends of the SNR timeseries due to filtering artifacts. The duration of the data stretch is chosen in a way that an in-band inspiral signal lies completely within a data stretch. The PSO algorithm explores each data stretch one after another. Data segments are downsampled till 1024 Hz as discussed later in Section III. PSDs used in the matched-filters are computed from the segments using Welch’s method (Welch, 1967).
@@
Based on the updated values of the fitness function of all the particles, the p-best and the g-best locations and fitnesses are computed as described in subsection II.2. Following this, Eq. 2-3 are used to assign a new set of positions to the particles for the next iteration. The algorithm also checks whether any particle has drifted beyond the desired search boundaries- a reflecting boundary puts back an outlier inside in a symmetric position while flipping its velocity 111While this is suitable in most cases, we find that an adsorbing, boundary is more effective for the eccentric dimension towards the recovery of non-eccentric sources, where instead of putting the outlier back to a symmetric location, we put it on the boundary.. The algorithm additionally checks for a terminal condition of a convergence below a small mean radius of the swarm or a preset timeout, whichever is earlier. The boundary conditions can improve the termination and also ensures no hang-up occurs due to non-availability of template parameters outside the region allowed by a waveform model. Note that these steps, namely evolving the position-velocity, applying the boundary conditions and the termination checks are computationally trivial. At this point, the current iteration is complete.
@@ II.4 Significance of candidates
When the iterations finish, the algorithm returns the detection template for a given stretch of data with the associated triggers from the individual detectors. The detection template point is already consistent in the source parameters across all the detectors. Now the secondary triggers are tested for any zero-lag time coincidence to create foreground events. A time coincidence test requires triggers from different detectors to occur within a time window defined by the light’s travel time between the detectors plus 5 ms accounting for any possible timing measurement errors. Several such foreground candidates, which pass both the parameter and the time coincidence tests, may still not be of astrophysical origin. To assess the possibility of a noise origin, one assigns a statistical significance to a foreground candidate and/or a probability of being of astrophysical origin, (Farr et al., 2015; Capano et al., 2017). Here we assign a false alarm rate (FAR) to the foreground events using the method of time-shifting of triggers obtained from processing the data immediately available near (or before) the candidate (Wąs et al., 2009). The algorithm maintains a sufficient buffer of the cumulative secondary triggers resulting from the detection templates of the neighbouring (or past) data stretches. Note that the different stretches can have their optimized g-best template points located throughout the parameter space. Thus, though the secondary triggers for a given data stretch have consistency over the source parameters, the collection of such triggers over multiple stretches do not enforce the source parameter consistency 222An appropriate strategy to utilize even the primary triggers for the background estimates in the PSO search can be explored in a future study. However, the cumulative triggers from the entire space are generally more prone to picking up abrupt noise and can even degrade the background statistic.. A test of accuracy of this FAR estimation strategy with real data is described in subsection III.3.
If the triggers obtained from a pair of detectors are artificially shifted with respect to each other by a time more than travel time of light between the two detectors, then the resulting coincidences cannot be of astrophysical origin. Such artificial background events are also ranked by their artificial network statistic value and a FAR estimate is assigned to a foreground candidate as
@@ III.3 Background event rates
In the previous sections, we have shown that the eccentric searches can outperform the aligned-spin searches in the presence of Gaussian noise. However, the data from a real detector can contain nearby noise glitches which can overwhelm the searches. To investigate this, we start with an hypothesis that there is no strong astrophysical signal and data segments contain only detector noise and glitches. We perform the searches independently using the same configuration as described in section III on several data blocks from the third observing run (O3a) of the Advanced LIGOs not containing any obvious CBC candidates (Abbott et al., 2023, 2021e). For simplicity we use data from only two detectors H1 and L1. Each 4096 seconds data block is divided into segments of 256 seconds with a 50% overlap, equally shared with its neighbouring segments and the analysis procedure described in Sec. II.3 is followed. Thus each PSO instance effectively searches 128 seconds of data from the available detectors. Note that the triggers crossing an SNR threshold undergo the same signal consistency tests and the network reweighted SNR is used as the ranking statistic. For comparison, we also conduct the aligned-spin non-eccentric search using PSO on the same data segments with identical set-up. The results are summarized in Fig. 11.
@@
However, to assess the accuracy of the estimation of the false alarm rates described earlier in subsection II.4, we use the triggers obtained from analyzing the above datasets. The cumulative number of all the foreground candidates are plotted against a given inverse false alarm rate (IFAR) threshold in Fig. 12. Here the FARs are assigned using the combined triggers from both BNS and BBH searches, inclusive of those forming the foreground candidates. The expected number of noise events for a given FAR threshold over a coincident observation time is calculated using Eq. 6. The figure demonstrates a general self consistency of the FARs assigned to any of the foreground candidates and indicates that the candidates are consistent with noise origin in this case.